OpenStudy (danise):

Write an explicit formula for the sequence and find the eighth term ? 1,2,4,8 PLEASE HELP! Algebra

OpenStudy (anonymous):

\[t _{n=}2^{n-1}\]

OpenStudy (anonymous):

put n=8 and simplify

OpenStudy (danise):

Can you help me find the eighth term?

OpenStudy (danise):

Simply the eight ?

OpenStudy (anonymous):

\[t1=1=2^0=2^{1-1}\] \[t2=2=2^1=2^{2-1}\] \[t3=4=2^2=2^{3-1}\] \[t4=8=2^3=2^{4-1}\] \[t _{8}=2^{8-1}=2^7=2*2*2*2*2*2*2=128\]